# Optimus-1[Uncaptioned image]: Hybrid Multimodal Memory Empowered Agents Excel in Long-Horizon Tasks

### v1

@@ Abstract
Building a general-purpose agent is a long-standing vision in the field of artificial intelligence. Existing agents have made remarkable progress in many domains, yet they still struggle to complete long-horizon tasks in an open world. We attribute this to the lack of necessary world knowledge and multimodal experience that can guide agents through a variety of long-horizon tasks. In this paper, we propose a Hybrid Multimodal Memory module to address the above challenges. It 1) transforms knowledge into Hierarchical Directed Knowledge Graph that allows agents to explicitly represent and learn world knowledge, and 2) summarises historical information into Abstracted Multimodal Experience Pool that provide agents with rich references for in-context learning. On top of the Hybrid Multimodal Memory module, a multimodal agent, Optimus-1, is constructed with dedicated Knowledge-guided Planner and Experience-Driven Reflector, contributing to a better planning and reflection in the face of long-horizon tasks in Minecraft. Extensive experimental results show that Optimus-1 significantly outperforms all existing agents on challenging long-horizon task benchmarks, and exhibits near human-level performance on many tasks. In addition, we introduce various Multimodal Large Language Models (MLLMs) as the backbone of Optimus-1. Experimental results show that Optimus-1 exhibits strong generalization with the help of the Hybrid Multimodal Memory module, outperforming the GPT-4V baseline on many tasks. Please see the project page at https://cybertronagent.github.io/Optimus-1.github.io/.
@@ 1 Introduction
Optimus Prime faces complex tasks alongside humans in Transformers to protect the peace of the planet. Creating an agent like Optimus that can perceive, plan, reflect, and complete long-horizon tasks in an open world has been a longstanding aspiration in the field of artificial intelligence [31, 15, 12]. Early research [1, 6, 17] developed simple agents by constructing policy networks. Recent works [34, 39, 37] have utilized Large Language Models (LLMs) as action planners for agents, generating executable sub-goal sequences for low-level action controllers. Leveraging the powerful instruction-following and logical reasoning capabilities of (Multimodal) LLMs, LLM-based agents have achieved remarkable success across multiple domains [13, 8, 9, 42]. Nevertheless, the ability of these agents to complete long-horizon tasks still falls significantly short of human-level performance.
@@
Insufficient Exploration of Structured Knowledge: Structured knowledge, encompassing open world rules, object relationships, and interaction methods with the environment, is essential for agents to complete complex tasks [25, 31]. However, existing agents [1, 17, 6] only learn dispersed knowledge from video data and are unable to efficiently represent and learn this structured knowledge, rendering them incapable of performing complex tasks.
@@ 2 Optimus-1
To build an agent that can perform long-horizon tasks in an open world like a human, we develop a multimodal composable agent, Optimus-1, in Minecraft. Optimus-1 consists of Hybrid Multimodal Memory, Knowledge-Guided Planner, Experience-Driven Reflector, and Action Controller.
@@ 2.1.1 Abstracted Multimodal Experience Pool
Relevant studies [4, 20, 16, 14] highlight the importance of historical information for agents completing long-horizon tasks. Minedojo [6] and Voyager [34] employed unimodal storage of historical information. Jarvis-1 [39] used a multimodal experience mechanism, storing task planning and visual information without summarization, leading to storage and retrieval challenges. To address this, we propose AMEP, which aims to summarize all multimodal information during the agent’s execution of the task. It can maintain long sequential data integrity, and improve storage and retrieval efficiency.
@@
Moreover, multimodal experience are also stored as failure cases when the feedback from the reflection phase is REPLAN. Thus, during the reflection phase, Optimus-1 can retrieve the most relevant cases from each of the three scenarios COMPLETE, CONTINUE, and REPLAN from AMEP as references. Experimental Results in Section 3.3 demonstrate the effectiveness of this innovative method of incorporating failure cases into in-context learning.
@@ 2.1.2 Hierarchical Directed Knowledge Graph
World knowledge is essential for an agent’s ability to perform long-horizon complex tasks. Instead of implicit learning through fine-tuning [24, 45], we propose HDKG, enabling agents to learn explicitly by retrieving knowledge graph.
Minecraft’s mining and crafting represent a complex knowledge network crucial for effective task planning. For instance, crafting a diamond sword [Uncaptioned image] requires two diamonds [Uncaptioned image] and one wooden stick [Uncaptioned image], while mining diamonds requires an iron pickaxe [Uncaptioned image], which involving further materials and steps. To enable a more structured knowledge representation, as shown in the Figure 2, we transform this knowledge into a graph $\mathcal{D}(\mathcal{V},\mathcal{E})$, where nodes set $\mathcal{V}$ represent objects, and directed edges set $\mathcal{E}$ point to nodes that can be crafted by this object. An edge $e\in\mathcal{E}$ in the $\mathcal{D}$ can be represented as $e=(u,v)$, where $u,v\in\mathcal{V}$. The directed graph efficiently stores and updates knowledge. For a given object $x$, retrieving the corresponding node allows extraction of a sub-graph $\mathcal{D}_{j}(\mathcal{V}_{j},\mathcal{E}_{j})\in\mathcal{D}$, where nodes set $\mathcal{V}_{j}$ and edges set $\mathcal{E}_{j}$ can be formulated as:
@@ 3.1 Experiments Setting
Environment. To ensure realistic gameplay like human players, we employ MineRL [10] with Minecraft 1.16.5 as our simulation environment. The agent operates at a fixed speed of 20 frames per second and only interacts with the environment via low-level action control signals of the mouse and keyboard. For more information about the detailed descriptions of the observation and action spaces, please refer to the Appendix C.
Table 1: Main Result of Optimus-1 on long-horizon tasks benchmark. We report the average success rate (SR), average number of steps (AS), and average time (AT) on each task group, the results of each task can be found in the Appendix G. Lower AS and AT metrics mean that the agent is more efficient at completing the task, while +$\infty$+ ∞ indicates that the agent is unable to complete the task. Overall represents the average result on the five groups of Iron, Gold, Diamond, Redstone, and Armor.
@@
Benchmark. We constructed a benchmark of 67 tasks to evaluate the Optimus-1’s ability to complete long-horizon tasks. As illustrated in Table 5, we divide the 67 Minecraft tasks into 7 groups according to recommended categories in Minecraft. Please refer to Appendix E for more details.
Baseline. We compare Optimus-1 with various agents, including GPT-3.5 https://openai.com/research/gpt-3.5, GPT-4V, DEPS [38], and Jarvis-1 [39] on the challenging long-horizon tasks benchmark. In addition, we employed 10 volunteers to perform the same task on the benchmark, and their average performance served as a human-level baseline. Please refer to Appendix E.2 for more details about human-level baseline. For a more comprehensive comparison, we also report Optimus-1’s performances on the benchmark used by Voyager [34], MP5 [24], and DEPS [38] in the Appendix G.2. Note that we initialize Optimus-1 with an empty inventory, while DEPS [38] and Jarvis-1 [39] have tools in their initial state. This makes it more challenging for Optimus-1 to perform the same tasks.
3.2 Experimental Results
[…]
@@ 3.4 Generalization Ability
(a) Generalisation of Hybrid Multimodal Memory
(b) Self-Evolution
@@ 3.5 Self-Evolution via Hybrid Multimodal Memory
As shown in Section 2.3, we randomly initialize the Hybrid Multimodal Memory of Optimus-1, then update it multiple times by using the "free exploration-teacher guidance" learning method. We set the epoch to 4, and the number of learning tasks to 160. At each period, Optimus-1 performs free exploration on 150 tasks and teacher guidance learning on the remaining 10 tasks, we then evaluate Optimus-1’s learning ability on the task groups same as ablation study. Experimental results are shown in Figure 5. It reveals that Optimus-1 keeps getting stronger through the continuous expansion of memory during the learning process of multiple periods. Moreover, it demonstrates that MLLM with Hybrid Multimodal Memory can incarnate an expert agent in a self-evolution manner [32].
@@ 4 Related Work
Agents in Minecraft. We summarise the differences of existing Minecraft agents in the Appendix E.3. Earlier work [21, 43, 2, 3] introduced policy models for agents to perform simple tasks in Minecraft. MineCLIP [6] used text-video data to train a contrastive video-language model as a reward model for policy, while VPT [1] pre-trained on unlabelled videos but lacked instruction as input. Building on VPT and MineCLIP, STEVE-1 [17] added text input to generate low-level action sequences from human instructions and images. However, these agents struggle with complex tasks due to limitations in instruction comprehension and planning. Recent work [37, 34, 46] incorporated LLMs as planning and reflection modules, but lacked visual information integration for adaptive planning. MP5 [24], MineDreamer [45], and Jarvis-1 [39] enhanced situation-aware planning by obtaining textual descriptions of visual information, yet lacked detailed visual data. Optimus-1 addresses these issues by directly using observation as situation-aware conditions in the planning phase, enabling more rational, visually informed planning. Additionally, unlike other agents requiring multiple queries for task refinement, Optimus-1 generates a complete and effective plan in one step with the help of HDKG. This makes Optimus-1 planning more efficient.
Memory in Agents. In the agent-environment interaction process, memory is key to achieving experience accumulation, environment exploration, and knowledge abstraction [44]. There are two forms to represent memory content in LLM-based agents: textual form [16, 14, 22] and parametric form [4, 20, 35]. In textual form, the information is explicitly retained and recalled by natural languages. In parametric form, the memory information is encoded into parameters and implicitly influences the agent’s actions. Recent work [36, 40, 11] has explored the long-term visual information storage and summarisation in MLLM. Our proposed hybrid multimodal memory module is plug-and-play and can provide world knowledge and multimodal experience for Optimus-1 efficiently.
@@ Appendix A Broader Impact
Appendix B Limitation and Future Work
In the framework of Optimus-1, we are dedicated to leverage proposed Hierarchical Directed Knowledge Graph and Abstracted Multimodal Experience Pool can be used to enhance the agent’s ability to plan and reflect. For Action Controller, we directly introduce STEVE-1 [17] as a generator of low-level actions. However, limited by STEVE-1’s ability to follow instructions and execute complex actions, Optimus-1 is weak in completing challenging tasks such as "beat ender dragon" and "build a house".
In order to further train STEVE-1’s low-level action generation ability, high-quality video-action data is indispensable. With the help of our highly scalable and adaptable architecture, we can obtain such high quality video data during inference process of Optimus-1. In our future work, we will strengthen the ability of Action Controller through further training.
Appendix C Minecraft
Minecraft is an extremely popular sandbox video game developed by Mojang Studios https://www.minecraft.net/en-us/article/meet-mojang-studios. It allows players to explore a blockly, procedurally generated 3D world with infinite terrain, discover and extract raw materials, craft tools and items, and build structures or earthworks(shown in Figure 6). In this game, AI agents need to face situations that are highly similar to the real world, making judgments and decisions to deal with various environments and problems. Therefore, Minecraft is a very suitable environment to be used as an AI testing simulator.
@@ Appendix C Minecraft
C.1 Basic Rules
Biomes. The Minecraft world is divided into different areas called "biomes". Different biomes contain different blocks and plants and change how the land is shaped. There are 79 biomes in Minecraft 1.16.5, including ocean, plains, forest, desert, etc. Diverse environments have high requirements for the generalization of agents.
@@ C.1 Basic Rules
C.2 Observation and Action Spaces
@@ 14 Craft - Execute a crafting recipe to obtain new item
C.3 Long-horizon Tasks
@@
For example, the agent now wants to complete the task "Craft iron sword [Uncaptioned image]". The agent needs to craft wood-level tools to mine stone [Uncaptioned image], and craft stone-level tools to mine iron ore [Uncaptioned image]. In order to craft tools, the agent needs a crafting table [Uncaptioned image]. To smelt iron ore [Uncaptioned image] into iron ingot [Uncaptioned image], the agent needs a furnace [Uncaptioned image]. Moreover, craft crafting table needs $4$ planks, and craft furnace needs $8$ cobblestone. In summary, the agent needs to obtain many raw materials, wood-level and stone-level tools, $1$ crafting table, $1$ furnace, and most importantly, $2$ iron ingots. The process of this task is shown in Figure 7.

### v2

@@ Abstract
Building a general-purpose agent is a long-standing vision in the field of artificial intelligence. Existing agents have made remarkable progress in many domains, yet they still struggle to complete long-horizon tasks in an open world. We attribute this to the lack of necessary world knowledge and multimodal experience that can guide agents through a variety of long-horizon tasks. In this paper, we propose a Hybrid Multimodal Memory module to address the above challenges. It 1) transforms knowledge into Hierarchical Directed Knowledge Graph that allows agents to explicitly represent and learn world knowledge, and 2) summarises historical information into Abstracted Multimodal Experience Pool that provide agents with rich references for in-context learning. On top of the Hybrid Multimodal Memory module, a multimodal agent, Optimus-1, is constructed with dedicated Knowledge-guided Planner and Experience-Driven Reflector, contributing to a better planning and reflection in the face of long-horizon tasks in Minecraft. Extensive experimental results show that Optimus-1 significantly outperforms all existing agents on challenging long-horizon task benchmarks, and exhibits near human-level performance on many tasks. In addition, we introduce various Multimodal Large Language Models (MLLMs) as the backbone of Optimus-1. Experimental results show that Optimus-1 exhibits strong generalization with the help of the Hybrid Multimodal Memory module, outperforming the GPT-4V baseline on various tasks. Please see the project page at https://cybertronagent.github.io/Optimus-1.github.io/.
@@ 1 Introduction
Optimus Prime faces complex tasks alongside humans in Transformers to protect the peace of the planet. Creating an agent [43, 13] like Optimus that can perceive, plan, reflect, and complete long-horizon tasks in an open world has been a longstanding aspiration in the field of artificial intelligence [22, 35, 36]. Early research developed simple policy through reinforcement learning [7] or imitation learning [1, 25]. A lot of work [46, 49] have utilized Large Language Models (LLMs) as action planners for agents, generating executable sub-goal sequences for low-level action controllers. Further, recent studies [51, 32] employed Multimodal Large Language Models (MLLMs) [4, 38, 55] as planner and reflector. Leveraging the powerful instruction-following and logical reasoning capabilities of (Multimodal) LLMs [24], LLM-based agents have achieved remarkable success across multiple domains [14, 9, 10, 54]. Nevertheless, the ability of these agents to complete long-horizon tasks still falls significantly short of human-level performance.
@@
Insufficient Exploration of Structured Knowledge: Structured knowledge, encompassing open world rules, object relationships, and interaction methods with the environment, is essential for agents to complete complex tasks [33, 43]. However, MLLMs such as GPT-4V ***https://openai.com/index/gpt-4v-system-card/ lack sufficient knowledge in Minecraft. Existing agents [1, 25, 7] only learn dispersed knowledge from video data and are unable to efficiently represent and learn this structured knowledge, rendering them incapable of performing complex tasks.
@@ 2 Optimus-1
In this section, we first elaborate on how to implement the Hybrid Multimodal Memory in Sec 2.1. As a core innovation, it plays a crucial role in enabling Optimus-1 to execute long-horizon tasks. Next, we give an overview of Optimus-1 framework (Sec 2.2), which consists of Hybrid Multimodal Memory, Knowledge-Guided Planner, Experience-Driven Reflector, and Action Controller. Finally, we introduce a non-parametric learning approach to expand the hybrid multimodal memory (Sec 2.3), thereby enhancing the success rate of task execution for Optimus-1.
@@ 2.1.1 Abstracted Multimodal Experience Pool
Relevant studies [23, 28, 17, 15] highlight the importance of historical information for agents completing long-horizon tasks. Minedojo [7] and Voyager [46] employed unimodal storage of historical information. Jarvis-1 [51] used a multimodal experience mechanism that stores task planning and visual information without summarization, posing challenges to storage capacity and retrieval speed. To address this issue, we propose AMEP, which aims to dynamically summarize all multimodal information during task execution. It preserves the integrity of long-horizon data while enhancing storage and retrieval efficiency.
@@ 2.1.2 Hierarchical Directed Knowledge Graph
In Minecraft, mining and crafting represent a complex knowledge network crucial for effective task planning. For instance, crafting a diamond sword [Uncaptioned image] requires two diamonds [Uncaptioned image] and one wooden stick [Uncaptioned image], while mining diamonds requires an iron pickaxe [Uncaptioned image], which involving further materials and steps. Such knowledge is essential for an agent’s ability to perform long-horizon complex tasks. Instead of implicit learning through fine-tuning [32, 58], we propose HDKG, which transforms knowledge into a graph representation. It enables the agent to perform explicit learning by retrieving information from the knowledge graph.
As shown in the Figure 2, we transform knowledge into a graph $\mathcal{D}(\mathcal{V},\mathcal{E})$, where nodes set $\mathcal{V}$ represent objects, and directed edges set $\mathcal{E}$ point to nodes that can be crafted by this object. An edge $e\in\mathcal{E}$ in the $\mathcal{D}$ can be represented as $e=(u,v)$, where $u,v\in\mathcal{V}$. The directed graph efficiently stores and updates knowledge. For a given object $x$, retrieving the corresponding node allows extraction of a sub-graph $\mathcal{D}_{j}(\mathcal{V}_{j},\mathcal{E}_{j})\in\mathcal{D}$, where nodes set $\mathcal{V}_{j}$ and edges set $\mathcal{E}_{j}$ can be formulated as:
@@ 2.2 Optimus-1: Framework
During task execution, even in cases where task failure necessitates REPLAN, multimodal experiences are stored in AMEP. Thus, during the reflection phase, Optimus-1 can retrieve the most relevant cases from each of the three scenarios COMPLETE, CONTINUE, and REPLAN from AMEP as references. Experimental Results in Section 3.3 demonstrate the effectiveness of this innovative method of incorporating failure cases into in-context learning.
@@ 3.1 Experiments Setting
Environment. To ensure realistic gameplay like human players, we employ MineRL [11] with Minecraft 1.16.5 as our simulation environment. The agent operates at a fixed speed of 20 frames per second and only interacts with the environment via low-level action control signals of the mouse and keyboard. For more information about the detailed descriptions of the observation and action spaces, please refer to the Appendix B.
Table 1: Main Result of Optimus-1 on long-horizon tasks benchmark. We report the average success rate (SR), average number of steps (AS), and average time (AT) on each task group, the results of each task can be found in the Appendix F. Lower AS and AT metrics mean that the agent is more efficient at completing the task, while +$\infty$+ ∞ indicates that the agent is unable to complete the task. Overall represents the average result on the five groups of Iron, Gold, Diamond, Redstone, and Armor.
@@
Benchmark. We constructed a benchmark of 67 tasks to evaluate the Optimus-1’s ability to complete long-horizon tasks. As illustrated in Table 5, we divide the 67 Minecraft tasks into 7 groups according to recommended categories in Minecraft. Please refer to Appendix D for more details.
Baseline. We compare Optimus-1 with various agents, including GPT-3.5  [^2] https://openai.com/research/gpt-3.5, GPT-4V, DEPS [50], and Jarvis-1 [51] on the challenging long-horizon tasks benchmark. In addition, we employed 10 volunteers to perform the same task on the benchmark, and their average performance served as a human-level baseline. Please refer to Appendix D.2 for more details about human-level baseline. For a more comprehensive comparison, we also report Optimus-1’s performances on the benchmark used by Voyager [46], MP5 [32], and DEPS [50] in the Appendix F.2. Note that we initialize Optimus-1 with an empty inventory, while DEPS [50] and Jarvis-1 [51] have tools in their initial state. This makes it more challenging for Optimus-1 to perform the same tasks.
@@
3.2 Experimental Results
[…]
(a) Generalisation of Hybrid Multimodal Memory
(b) Self-Evolution
@@ 3.5 Self-Evolution via Hybrid Multimodal Memory
As shown in Section 2.3, we randomly initialize the Hybrid Multimodal Memory of Optimus-1, then update it multiple times by using the “free exploration-teacher guidance” learning method. We set the epoch to 4, and the number of learning tasks to 160. At each period, Optimus-1 performs free exploration on 150 tasks and teacher guidance learning on the remaining 10 tasks, we then evaluate Optimus-1’s learning ability on the task groups same as ablation study. Experimental results are shown in Figure 5. It reveals that Optimus-1 keeps getting stronger through the continuous expansion of memory during the learning process of multiple periods. Moreover, it demonstrates that MLLM with Hybrid Multimodal Memory can incarnate an expert agent in a self-evolution manner [44].
@@ 4 Related Work
4.1 Agents in Minecraft
We summarise the differences of existing Minecraft agents in the Appendix D.3. Earlier work [29, 56, 2, 3] introduced policy models for agents to perform simple tasks in Minecraft. MineCLIP [7] used text-video data to train a contrastive video-language model as a reward model for policy, while VPT [1] pre-trained on unlabelled videos but lacked instruction as input. Building on VPT and MineCLIP, STEVE-1 [25] added text input to generate low-level action sequences from human instructions and images. However, these agents struggle with complex tasks due to limitations in instruction comprehension and planning. Recent work [49, 46, 59] incorporated LLMs as planning and reflection modules, but lacked visual information integration for adaptive planning. MP5 [32], MineDreamer [58], and Jarvis-1 [51] enhanced situation-aware planning by obtaining textual descriptions of visual information, yet lacked detailed visual data. Optimus-1 addresses these issues by directly using observation as situation-aware conditions in the planning phase, enabling more rational, visually informed planning. Additionally, unlike other agents requiring multiple queries for task refinement, Optimus-1 generates a complete and effective plan in one step with the help of HDKG. This makes Optimus-1 planning more efficient.
4.2 Memory in Agents
In the agent-environment interaction process, memory is key to achieving experience accumulation [21], environment exploration [16], and knowledge abstraction [57]. There are two forms to represent memory content in LLM-based agents: textual form [17, 15, 30] and parametric form [5, 28, 47, 20]. In textual form, the information is explicitly retained and recalled by natural languages. In parametric form, the memory information [37] is encoded into parameters and implicitly influences the agent’s actions. Recent work [48, 52, 12] has explored the long-term visual information storage [18, 19] and summarisation in MLLM. Our proposed hybrid multimodal memory module is plug-and-play and can provide world knowledge and multimodal experience for Optimus-1 efficiently.
@@ 5 Conclusion
6 Limitation and Future Work
In the framework of Optimus-1, we are dedicated to leverage proposed Hierarchical Directed Knowledge Graph and Abstracted Multimodal Experience Pool can be used to enhance the agent’s ability to plan and reflect. For Action Controller, we directly introduce STEVE-1 [25] as a generator of low-level actions. However, limited by STEVE-1’s ability to follow instructions and execute complex actions, Optimus-1 is weak in completing challenging tasks such as “beat ender dragon” and “build a house”. Therefore, a potential future research direction is to enhance the instruction following and action generation capabilities of action controller.
In addition, most of the work, including Optimus-1, utilize a multimodal large language model for planning and reflection, which then drives an action controller to perform the task. Building an end-to-end vision-language-action agent will be future work.
7 Acknowledgement
This study is supported by National Natural Science Foundation of China (Grant No. 62236003 and 62306090), Shenzhen College Stability Support Plan (Grant No. GXWD20220817144428005), Natural Science Foundation of Guangdong Province of China (Grant No. 2024A1515010147), and Major Key Project of Peng Cheng Laboratory (Grant No. PCL2023A08).
@@ Appendix A Broader Impact
Appendix B Minecraft
Minecraft is an extremely popular sandbox video game developed by Mojang Studios  [^3] https://www.minecraft.net/en-us/article/meet-mojang-studios. It allows players to explore a blockly, procedurally generated 3D world with infinite terrain, discover and extract raw materials, craft tools and items, and build structures or earthworks (shown in Figure 6). Minecraft is a valuable and representative environment for evaluating long-horizon tasks, offering greater diversity and complexity compared to other environments. Unlike web/app navigation [54] and embodied manipulation [16], Minecraft is an open world with a complex and dynamic environment (79 biomes, including ocean, plains, forest, desert, etc.). To complete long-horizon tasks, agents must achieve multiple sub-goals (e.g., 15 sub-goals to craft a diamond sword), making the construction of a Minecraft agent quite challenging. Many studies [46, 32, 51] have chosen Minecraft as the environment for validating performance on long-horizon tasks. Extensive experimental results in the paper show that Optimus-1 outperforms all baselines. Therefore, we chose Minecraft as open-world environment to evaluate the ability of agents to perform long-horizon tasks.
@@ Appendix B Minecraft
B.1 Basic Rules
Biomes. The Minecraft world is divided into different areas called “biomes”. Different biomes contain different blocks and plants and change how the land is shaped. There are 79 biomes in Minecraft 1.16.5, including ocean, plains, forest, desert, etc. Diverse environments have high requirements for the generalization of agents.
@@ B.1 Basic Rules
B.2 Observation and Action Spaces
@@ 14 Craft - Execute a crafting recipe to obtain new item
B.3 Long-horizon Tasks
@@
For example, the agent now wants to complete the task “Craft iron sword [Uncaptioned image]”. The agent needs to craft wood-level tools to mine stone [Uncaptioned image], and craft stone-level tools to mine iron ore [Uncaptioned image]. In order to craft tools, the agent needs a crafting table [Uncaptioned image]. To smelt iron ore [Uncaptioned image] into iron ingot [Uncaptioned image], the agent needs a furnace [Uncaptioned image]. Moreover, craft crafting table needs $4$ planks, and craft furnace needs $8$ cobblestone. In summary, the agent needs to obtain many raw materials, wood-level and stone-level tools, $1$ crafting table, $1$ furnace, and most importantly, $2$ iron ingots. The process of this task is shown in Figure 7.